OpenStudy (diamond77108):

plz help 2/3+y−4/5=1/3

OpenStudy (leenathan):

hello and welcome to open study for further assistance please go to brainly.com

OpenStudy (otherworldly):

y = -2/3 + 1/3 + 4/5

OpenStudy (otherworldly):

make the denominator the same in all the fractions

OpenStudy (otherworldly):

y = -10/15 + 5/15 + 12/15

OpenStudy (otherworldly):

y = 7/15

OpenStudy (otherworldly):

good luck *thumbs up* :)

OpenStudy (diamond77108):

thx

OpenStudy (diamond77108):

i kept thinking it was -7/15

OpenStudy (otherworldly):

no it wouldn't be because -10 +5 + 12 is 7 ........-10 +12 = 2 + 5 which is 7 :)